# QTabGAN: A Hybrid Quantum-Classical GAN for Tabular Data Synthesis

### v1

@@ 2.5 Generative Adversarial Networks (GANs)
Generative Adversarial Networks are a class of generative models that learn the underlying distribution of a dataset through adversarial training. Figure 1 represents the basic architecture of Generative Adversarial Networks(GANs). A GAN consists of two adversarial components: the generator that maps an input source such as random noise, conditioning information (e.g., class labels), or both together into synthetic data samples, and the discriminator that evaluates whether the given sample is real or synthetic. The neural networks are trained using an adversarial minimax objective where the generator aims to produce samples that the discriminator cannot distinguish, while the discriminator aims to improve its ability to distinguish between the real and synthetic data. This adversarial training enables the generator to gradually synthesize high-fidelity samples that resemble the real-world data. The minimax objective function of a typical GAN is given as follows:
@@ 3 Model Architecture
Our proposed framework for QTabGAN is a hybrid quantum-classical framework. Figure 2 represents the model architecture of QTabGAN. The generator’s core is a Variational Quantum Circuit (VQC) that generates the probability distributions for synthetic data. VQCs are used to exploit quantum superposition and entanglement. This enables the generator to capture complex data distributions and correlations that classical generators struggle to capture. Next, the data distributions are mapped to the features of the tabular data using a classical neural network. Finally, the classical discriminator distinguishes between real and synthetic data.
@@ 3.2 Discriminator Architecture
The discriminator is a densely connected classical feedforward neural network designed to distinguish real tabular samples from synthetic samples generated by the hybrid quantum-classical generator during adversarial training. Figure 4 illustrates the discriminator architecture. The discriminator provides feedback to the generator, which guides the generator to produce synthetic samples that closely match the real data distribution.
@@ 4 Experimental Setup
Figure 5: Evaluation flow for ML Utility of Classification datasets
@@ 5.1 Dataset Description
To evaluate our QTabGAN, we have used seven different datasets in total. The dataset descriptions are given below. The first two datasets are used for regression, while the remaining five are used for classification.
@@ 5.3 Result Analysis
Figure 6 and Figure 7 show the results for the classification and the regression datasets, respectively. For both classification and regression datasets, we compare QTabGAN against eight state-of-the-art generative adversarial network (GAN) models, namely, CTAB-GAN+, CTAB-GAN, IT-GAN, TVAE, CTGAN, TableGAN, MedGAN, and CW-GAN, to evaluate their relative performance across multiple metrics. Among these baselines, CTAB-GAN+ [34] is used as the main state-of-the-art baseline for comparison. We also compare QTabGAN against the TabularQGAN [7] model, which, to our knowledge, is the only other published work on tabular data generation using quantum GANs. For the comparison of ML utility metrics, QTabGAN is evaluated using the full set of features available in each dataset, whereas TabularQGAN is evaluated using a reduced subset of three to four features, as the TabularQGAN architecture does not scale to higher-dimensional feature spaces. In contrast, for the statistical similarity evaluation, both QTabGAN and TabularQGAN are assessed using the same subset of features to ensure a fair and consistent comparison, since these metrics are inherently feature-dependent whereas ML utility metrics does not depend upon number of features. For both classification and regression datasets, the results for each metric are obtained by averaging the corresponding metric values across all datasets. We first discuss the results on ML utility, followed by an analysis of statistical similarity.
@@ 5.3.1 ML Utility
A key limitation of the TabularQGAN model is that, for the same number of qubits, it can model only a small subset of features (typically three to four features), whereas QTabGAN is able to generate all features. This limitation directly affects the downstream performance of TabularQGAN, which can be seen in its consistently lower ML utility metrics across both classification and regression datasets. Furthermore, we can also observe that TabularQGAN does not even outperform classical models across the ML utility metrics.
From the results above, it can be seen that QTabGAN outperforms all other state-of-the-art methods across all ML utility metrics for both classification and regression datasets.
@@ 5.3.2 Statistical Similarity
We also compare QTabGAN with the TabularQGAN model using statistical similarity metrics across both classification and regression datasets. To maintain a fair comparison in terms of statistical similarity, we evaluate similarity scores on the same features used by the TabularQGAN model for all the datasets. Table 3 compares the statistical similarity between real and synthetic data generated by QTabGAN and TabularQGAN across both classification and regression datasets. Lower values indicate better alignment with the real data distributions.
@@
5.3.3 Best and Worst Case Analysis
Table 2 presents the accuracy and F1-score values obtained using real and synthetic data for the best-performing dataset (Adult) and one of the worst-performing datasets (Loan). For the Adult dataset, the differences in Accuracy and F1-score between models trained on real and synthetic data are the smallest among all classification datasets considered, indicating that QTabGAN’s synthetic samples closely match the underlying distribution of the real data.
@@ 5.3.4 CDF-Based Distribution Analysis
Figure 8 presents smoothed cumulative distribution function (CDF) curves comparing real and synthetic samples for best and worst features across a classification and a regression dataset. CDFs provide an intuitive way to characterize data distributions, and smoothing helps reveal overall trends while reducing noise, enabling a clearer assessment of how well the synthetic data reproduces the statistical properties of the real data.
Sub-figure 8(a) and Sub-figure 8(b) show the best and worst CDF plots of the Adult dataset for age and capital-gain features, respectively. The CDF for the age feature shows a strong overlap between the real and synthetic data, indicating that QTabGAN effectively captures the overall distribution. For capital-gain, small differences are visible in some regions; however, the synthetic CDF closely follows the general cumulative trend of the real data. Such behaviour is expected for highly skewed and heavy-tailed features like capital-gain. Overall, QTabGAN preserves the main distributional characteristics across both smoothly varying and more challenging features, demonstrating stable and robust distributional learning.
Sub-figure 8(c) and Sub-figure 8(d) show the best and worst plots of the King dataset for sqft_living and yr_built features, respectively. The CDF for sqft_living15 closely matches the real and synthetic data, indicating that QTabGAN effectively captures the cumulative distribution. For yr_built, the CDF shows a smooth and monotonic trend, indicating that the proposed model is able to learn the overall distribution of the feature across its full range. The close alignment of the curves suggests that both lower and higher values are adequately represented, reflecting stable learning behaviour. These results show that the model preserves the global statistical characteristics in both cases and exhibits consistent performance across features in the King dataset.
@@
Method Classification Datasets Regression Datasets
Avg. JSD Diff. Corr. Diff. Corr.
QTabGAN 0.05 0.4 0.14
TabularQGAN 0.20 0.6 0.19

### v2

@@ 2.5 Generative Adversarial Networks (GANs)
Generative Adversarial Networks are a class of generative models that learn the underlying distribution of a dataset through adversarial training. The basic architecture of Generative Adversarial Networks(GANs) is shown in Figure 1. A GAN consists of two adversarial components: the generator that maps an input source such as random noise, conditioning information (e.g., class labels), or both together into synthetic data samples, and the discriminator that evaluates whether the given sample is real or synthetic. The neural networks are trained using an adversarial minimax objective where the generator aims to produce samples that the discriminator cannot distinguish, while the discriminator aims to improve its ability to distinguish between the real and synthetic data. This adversarial training enables the generator to gradually synthesize high-fidelity samples that resemble the real-world data. The minimax objective function of a typical GAN is given as follows:
@@ 3 Model Architecture
Our proposed framework for QTabGAN is a hybrid quantum-classical framework. The model architecture of QTabGAN is shown in Figure 2. The generator’s core is a Variational Quantum Circuit (VQC) that generates the probability distributions for synthetic data. VQCs are used to exploit quantum superposition and entanglement. This enables the generator to capture complex data distributions and correlations that classical generators struggle to capture. Next, the data distributions are mapped to the features of the tabular data using a classical neural network. Finally, the classical discriminator distinguishes between real and synthetic data.
@@ 3.2 Discriminator Architecture
The discriminator is a densely connected classical feedforward neural network designed to distinguish real tabular samples from synthetic samples generated by the hybrid quantum-classical generator during adversarial training. The discriminator architecture is shown in Figure 4. The discriminator provides feedback to the generator, which guides the generator to produce synthetic samples that closely match the real data distribution.
@@ 4.2 QTabGAN Training
Figure 5: Evaluation flow for ML Utility of Classification datasets
@@ 5.1 Dataset Description
To evaluate our QTabGAN, we have used seven different datasets in total. The dataset descriptions are given below in Table 1. The first two datasets are used for regression, while the remaining five are used for classification.
@@ 5.3 Result Analysis
Next, we discuss the results obtained from our experiments on both classification and regression datasets. For both classification and regression datasets, we compare QTabGAN against eight state-of-the-art generative adversarial network (GAN) models, namely, CTAB-GAN+, CTAB-GAN, IT-GAN, TVAE, CTGAN, TableGAN, MedGAN, and CW-GAN, to evaluate their relative performance across multiple metrics. Among these baselines, CTAB-GAN+ [34] is used as the main state-of-the-art baseline for comparison. We also compare QTabGAN against the TabularQGAN [7] model, which, to our knowledge, is the only other published work on tabular data generation using quantum GANs. For both classification and regression datasets, the results for each metric are obtained by averaging the corresponding metric values across all datasets. The experimental results are illustrated in Figure 6 and Figure 7 for the classification and the regression datasets, respectively. We first discuss the results on ML utility, followed by an analysis of statistical similarity.
@@ 5.3.1 ML Utility
We observe that QTabGAN outperforms both classical GANs and TabularQGAN on downstream ML utility metrics. This behaviour is expected, as QTabGAN efficiently explores a much larger feature space than classical GANs do, thanks to VQC’s inherent computational capabilities. Moreover, QTabGAN’s better performance compared to TabularQGAN is attributed to their architectural differences. For a given number of qubits (say, ten qubits), TabularQGAN can model only a small number of features (typically three to four), whereas QTabGAN can model all the features by leveraging the classical mapper. This limitation directly affects TabularQGAN’s downstream performance, as evidenced by consistently lower ML utility metrics across both classification and regression datasets. Furthermore, we can also observe that TabularQGAN does not even outperform classical models across the ML utility metrics.
@@ 5.3.2 Statistical Similarity
We also compare QTabGAN with the TabularQGAN model using statistical similarity metrics across both classification and regression datasets. For the statistical similarity metrics evaluation, both QTabGAN and TabularQGAN are assessed using the same subset of features to ensure a fair and consistent comparison, since these metrics are inherently feature-dependent. The comparative results for both classification and regression datasets are summarized in Table 2. Lower values indicate better alignment with the real data distributions.
@@
Method Classification Datasets Regression Datasets
Avg. JSD Diff. Corr. Diff. Corr.
QTabGAN 0.05 0.4 0.14
TabularQGAN 0.20 0.6 0.19
5.3.3 Best and Worst Case Analysis
To better understand the model’s behaviour, we analyze two representative classification datasets: the best-performing dataset (Adult) and one of the worst-performing datasets (Loan). This comparison helps us better understand how QTabGAN behaves in both favourable and more complex scenarios. The detailed accuracy and F1-score results obtained using real and synthetic data for these datasets are summarized in Table 3. For the Adult dataset, the differences in Accuracy and F1-score between models trained on real and synthetic data are the smallest among all classification datasets considered, indicating that QTabGAN’s synthetic samples closely match the underlying distribution of the real data.
@@ 5.3.4 CDF-Based Distribution Analysis
To further examine the distributional alignment between real and synthetic data, we analyze representative features from both a classification dataset and a regression dataset. Specifically, we consider one of the best-matched features and one of the worst-matched features to understand how well the model captures different distribution patterns. The smoothed cumulative distribution function (CDF) curves for these features are shown in Figure 8. CDFs provide an intuitive way to characterize data distributions, and smoothing helps reveal overall trends while reducing noise, enabling a clearer assessment of how well the synthetic data reproduces the statistical properties of the real data.
For the Adult dataset, Sub-figure 8(a) and Sub-figure 8(b) show the best and worst CDF plots for age and capital-gain features, respectively. The CDF for the age feature shows a strong overlap between the real and synthetic data, indicating that QTabGAN effectively captures the overall distribution. For capital-gain, small differences are visible in some regions; however, the synthetic CDF closely follows the general cumulative trend of the real data. Such behaviour is expected for highly skewed and heavy-tailed features like capital-gain. This suggests that the model can effectively capture non-uniform and long-tailed distributions without introducing large deviations. Overall, QTabGAN preserves the main distributional characteristics across both smoothly varying and more challenging features, demonstrating stable and robust distributional learning.
For the King dataset, Sub-figure 8(c) and Sub-figure 8(d) show the best and worst plots for sqft_living and yr_built features, respectively. The CDF for sqft_living15 closely matches the real and synthetic data, indicating that QTabGAN effectively captures the cumulative distribution. For yr_built, the CDF shows a smooth and monotonic trend, indicating that the proposed model is able to learn the overall distribution of the feature across its full range. The close alignment of the curves suggests that both lower and higher values are adequately represented, reflecting stable learning behaviour. Importantly, the model maintains consistency not only in central regions of the distribution but also toward the tails, which are often more difficult to approximate accurately. These results show that the model preserves the global statistical characteristics in both cases and exhibits consistent performance across features in the King dataset.